# Holographic thermal propagator from modularity

### v1

@@ 2 Zamolodchikov’s $q$-recursion
$\mu_{1}=\lambda_{1}+\lambda_{2}\quad,\quad\mu_{2}=\lambda_{1}-\lambda_{2},\quad,\quad\mu_{3}=\lambda_{3}+\lambda_{4}\quad,\quad\mu_{4}=\lambda_{3}-\lambda_{4}\quad.$ (2.3)
@@
$R_{m,n}(\vec{\mu})=\frac{2\prod_{r,s}\prod_{i=1}^{4}\left(\mu_{i}-\lambda_{r,s}\right)}{\prod_{k,l}^{\prime}\lambda_{k,l}}$ (2.9)
@@ 4 $\mathcal{S}$-duality constraints on the prepotential: quasi-modular forms
$a_{D}=\tau\left(-a+\frac{\delta}{12a}\left(-h_{0}+a\frac{\partial F_{q}}{\partial a}\right)\right)$ (4.24)
@@
acts on $\tilde{\alpha}_{2n}/a^{2n}$. Due to (4.16) and the behaviour of $E_{2k}$ (see appendix A) this operation is equivalent to
$E_{2}\to E_{2}+\delta\quad,\quad a\to a-\frac{\delta}{12a}\left(-h_{0}+a\frac{\partial F_{q}}{\partial a}\right).$ (4.27)
@@
$=$ $h_{0}\log{\left(1-\frac{\delta}{12a^{2}}\left(-h_{0}+a\frac{\partial\tilde{H}}{\partial a}\right)\right)}-\frac{\delta}{24a^{2}}\left(-h_{0}+a\frac{\partial\tilde{H}}{\partial a}\right).$
@@
$\frac{\partial\tilde{H}}{\partial E_{2}}=\frac{1}{24a^{2}}\left(-h_{0}+a\frac{\tilde{H}}{\partial a}\right)^{2}.$ (4.29)

### v2

@@ 2 Zamolodchikov’s $q$-recursion
$\mu_{1}=\lambda_{1}+\lambda_{2}+\frac{Q}{2}\ ,\quad\mu_{2}=\lambda_{1}-\lambda_{2}+\frac{Q}{2}\ ,\quad\mu_{3}=\lambda_{3}+\lambda_{4}+\frac{Q}{2}\ ,\quad\mu_{4}=\lambda_{3}-\lambda_{4}+\frac{Q}{2}\ .$ (2.3)
@@
$R_{m,n}(\vec{\mu})=\frac{2\prod_{r,s}\prod_{i=1}^{4}\left(\mu_{i}-\frac{Q}{2}-\lambda_{r,s}\right)}{\prod_{k,l}^{\prime}\lambda_{k,l}}$ (2.9)
@@ 4 $\mathcal{S}$-duality constraints on the prepotential: quasi-modular forms
$a_{D}=\tau\left(-a+\frac{\delta}{12a}\left(-h_{0}+a\frac{\partial\tilde{H}}{\partial a}\right)\right)$ (4.24)
@@
acts on $\tilde{\alpha}_{2n}/a^{2n}$. Due to (4.16) and the behaviour of $E_{2k}$ (see appendix B) this operation is equivalent to
$E_{2}\to E_{2}+\delta\quad,\quad a\to a-\frac{\delta}{12a}\left(-h_{0}+a\frac{\partial\tilde{H}}{\partial a}\right).$ (4.27)
@@
$=$ $h_{0}\log{\left(1-\frac{\delta}{12a^{2}}\left(-h_{0}+a\frac{\partial\tilde{H}}{\partial a}\right)\right)}-\frac{\delta}{24a^{2}}\left(-h_{0}+a\frac{\partial\tilde{H}}{\partial a}\right)^{2}.$
@@
$\frac{\partial\tilde{H}}{\partial E_{2}}=\frac{1}{24a^{2}}\left(-h_{0}+a\frac{\partial\tilde{H}}{\partial a}\right)^{2}.$ (4.29)
@@ 5 Low temperature expansion of the black brane propagator
From Fig. 1, it appears that the Padé approximant and the numerical solution begin to deviate around $\hat{T}\approx 10^{-4}$. Although extending the convergence interval is beyond the scope of this paper, we verified that higher-order Padé approximants yield a better fit. In particular, convergence tends to be better for odd values of $n+m$. A rough estimate suggests that achieving a satisfactory fit up to $\hat{T}\approx 1$ would require $n+m\approx 1000$.